### 31 Jan 2024 v1

Abstract
Context:AD Leo is a young and active M dwarf with high flaring rates across the X-ray to radio bands. Flares accelerate particles in the outer coronal layers and often impact exo-space weather. Wide-band radio dynamic spectra let us explore the evolution of particle acceleration activity across the corona. Identifying the emission features and modelling the mechanisms can provide insights into the possible physical scenarios driving the particle acceleration processes.
Aims:We performed an 8 h monitoring of AD Leo across the 550 - 850 MHz band using upgraded-Giant Metrewave Radio Telescope (uGMRT). The possible flare and post-flare emission mechanisms are explored based on the evolution of flux density and polarisation.
Methods:A python based module, named VISAD, was developed to obtain the visibility averaged wide-band dynamic spectra. Direct imaging was also performed with different frequency-time averaging. Based on existing observational results on AD Leo and on solar active region models, radial profiles of electron density and magnetic fields were derived. Applying these models, we explore the possible emission mechanisms and magnetic field profile of the flaring active region.
Results:The star displayed high brightness temperature ( 10 – 10 K) throughout the observation. The emission was also nearly 100% left circularly polarised during bursts. Post flare phase is characterised by a highly polarised (60 - 80%) solar-like type IV burst confined above 700 MHz.
Conclusions:The flare emission favors a Z -mode or a higher harmonic X-mode electron cyclotron maser emission mechanism. The 700 MHz post-flare activity is consistent with a type-IV radio burst from flare-accelerated particles trapped in magnetic loops, which could be a coronal mass ejection (CME) signature. This is the first solar-like type-IV burst reported on a young active M dwarf belonging to a different age - related activity population (‘C’ branch) compared to the Sun (‘I’ branch). We also find that, a multipole expansion model of the active region magnetic field better accounts for the observed radio emission than a solar-like active region profile.
Energetic particle activity in AD Leo: detection of a solar-like type-IV burst
Key Words.:
stars: flare - stars: activity - stars: magnetic field - stars: coronae - radio continuum: stars - stars: low-mass1 Introduction
Low-mass main-sequence stars of spectral type M (M dwarfs) make up the most numerous stellar type in our galaxy. They also have the highest occurrence rate of Earth-like exoplanets in their habitable zones (Bashi et al., 2020). Meanwhile, M-dwarfs form the most active population among cool main-sequence stars with relatively high values of flaring rates, X-ray to bolometric flux ratio, and other activity metrics (see Donati & Landstreet, 2009, for a review). Moreover, the younger stars are more active with stronger magnetic fields and faster rotation rates (e.g., Vidotto et al., 2014; Lund et al., 2020). Stellar flares cause energetic space-weather events including the release of energetic particles, intense high-energy radiation and large-scale eruptions/outflows, which threaten the stability of planetary atmospheres and hence the habitability. So, it is important to understand the mechanism, evolution and space-weather consequences of M dwarf flares.
For this study, we chose the very active M3.5V star AD Leo located at 5 pc from the Sun. The star is proposed to host a planet (Tuomi et al., 2018), though the detection was later refuted (Carleo et al., 2020). The star has a typical mean optical flaring rate of 0.6 h (Dal, 2020). AD Leo displays high levels of chromospheric and coronal activity in several spectral lines and X-ray bands (Crespo-Chacón et al., 2006). Villadsen & Hallinan (2019) estimated a radio flaring rate of 0.2 – 0.3 h with a coherent burst emission duty cycle of 20 – 25% in the 0.34 – 1.4 GHz range.
Flare emission in the metrewave band is sensitive to various coherent emission signatures from flare-accelerated electrons either trapped in low-coronal magnetic field structures or streaming across coronal layers (e.g., Wild, 1970; McLean & Labrum, 1985; Bastian, 1990; Melrose, 2009, 2017a; Zic et al., 2020; Vedantham, 2021). The two prominant coherent emission mechanisms discussed in the literature include the electron cyclotron maser (ECME) emission and the plasma emission (see, Dulk, 1985; Güdel, 2002; Melrose, 2017a, for an overview). ECME dominates when the ratio of plasma frequency to gyro-frequency (/) at the flare source region is less than 1.5 (Melrose & Dulk, 1982; Melrose et al., 1984a; Dulk, 1985; Dulk & Winglee, 1987). When /3, the favoured emission mechanism is plasma emission (e.g., Ginzburg & Zhelezniakov, 1958; Tsytovich & Kaplan, 1969; Melrose & Sy, 1972; Cairns & Melrose, 1985; Melrose, 2009). Depending on the emission mechanism, the coherent emission frequencies correspond to either the fundamental mode or the harmonics of the local plasma frequency () or gyro-frequency (), letting us infer either the local density or magnetic field strength, respectively. Apart from the flare emission, M dwarfs are also known to display non-flaring (or non-impulsive) continuum emission with high brightness temperature () (e.g., White et al., 1989; Bastian, 1990). In a few GHz (¡2 GHz) bands, the continuum ranges from 10 - 10 K (Osten et al., 2008; Villadsen & Hallinan, 2019). In metrewavebands can be even higher (e.g., Bastian, 1994; Güdel, 2002; Villadsen & Hallinan, 2019). The continuum emission can form either from a gyrosynchrotron mechanism, or a combination of various coherent emission processes operating at multiple active regions (Bastian, 1996; Benz, 1996; Zaitsev et al., 1997).
There have been many studies that explored radio emission from AD Leo, reporting several flaring events. While instrumental limitations forced earlier studies to use only a few disjoint narrow frequency bands (e.g., Spangler et al., 1974; Bastian et al., 1990), recent observations with continuous wide-band instruments has enabled spectroscopic explorations revealing rich variability in the dynamic spectrum (e.g., Osten et al., 2008; Villadsen & Hallinan, 2019; Zhang et al., 2023). These studies have revealed several features like fast pulsations at scales of ms - min and high drift rates in the dynamic spectrum, often related to the propagation of energetic particles or spatio-temporal evolution of plasma instabilities in the associated active regions (e.g., Bastian et al., 1990; Osten & Bastian, 2006; Osten et al., 2008; Zhang et al., 2023). However, much of these observations had been close to or above 1 GHz. Here we present a 8 h long monitoring of AD Leo across a wide sub-GHz band, which is expected to probe higher atmospheric layers than those probed by GHz bands. Though, the original study was planned with simultaneous observation spanning from X-ray to radio band, only the radio observations were successful due to various technical issues. We present the results from meterwave observations here.
2 Data and Methodology
AD Leo was observed with uGMRT (Gupta et al., 2017) in Band 4 (550 - 850 MHz) from 3 December 2021, 21:17:41 UT to 4 December 2021, 04:17:41 UT at 5 s and 100 kHz time and frequency resolution, respectively. The data were gathered in RR and LL polarisations, letting us explore the spectro-temporal evolution of the circular polarisation along with brightness temperature () or flux density. The data analysis was done using the functionalities in the interferometric analysis package, Common Astronomy Software Applications (CASA; CASA Team et al., 2022). The data were first flagged and calibrated using standard procedures. The flux calibrator was 3C147, and the phase calibrator was 1021+219. Portions of data affected by radio frequency interference from terrestrial sources and systematic instrumental effects were identified and flagged manually. The automatic flagging routine in CASA, namely tfcrop, was also used. This was followed by bandpass calibration using the flux calibrator, and then by time dependent gain calibration using the phase calibrator. The bandpass solution and the time dependent gain solution was then extrapolated to the target star field. The calibrated target field data was further flagged using the Rflag routine in CASA. Initial imaging was done using the CASA task tclean in mtmfs mode, with nterms set to 2 averaging over the entire observing time. This produced a mean model image for the entire spectral band. This was taken as the initial model and self-calibration was performed in an iterative manner (Cornwell & Fomalont, 1999) to improve this model. After the second iteration, an optimal model was obtained. So we stopped at the second iteration and adopted the CLEAN model of the iteration as the mean model image of the field of view for the observation period. The AD Leo emission at the phase centre was masked from the final model image and the model visibilities were subtracted from the corrected data. The remaining corrected visibilities pertain to purely stellar emission. This stellar data was subjected to two forms of analysis as described below.
First, we developed a python based package, Visibility Averaged Dynamic spectrum (VISAD), to extract the visibilities from a calibrated CASA measurement set (MS) and generate a dynamic spectrum (DS) with required time and frequency averaging. The package has routines to perform additional data flagging on the calibrated MS, if required. The package uses CASA tasks inside a python environment and employs the parallel processing tools in python (multiprocessing111https://docs.python.org/3/library/multiprocessing.html#reference) enabling analysis of several hours of wide bandwidth datasets within a couple of minutes depending on the available machine cores. The module is made publicly available via GitHub222https://github.com/atul3790/Visibility-averaged-DS. We obtained both dynamic spectra and band averaged light curves in Stokes I and V using VISAD. The Stokes V data were corrected as per the suggestions in Das et al. (2020) to adhere to the standard IAU/IEEE convention.
We have also produced image spectral cubes for the entire observing period with varying time and frequency averaging, carefully chosen to detect the star with at least a signal to noise ratio of 5 in every spectro-temporal slice. This exercise generated two data sets: a library of stellar spectra for different time bins spanning the entire observation period and a band averaged stellar light curve. The spectra were generated with a resolution of 20 MHz and a time averaging of 2 - 3 min during the very active periods. During periods of weak activity a 50 MHz averaging with 5 - 15 min time averaging was employed. We used the image cube information to cross-verify the flux densities obtained from visibility averaged DS for the flare period. The values matched within 15 - 20%. A source of this discrepancy is the additional flagging done in a couple of higher frequency channels while using VISAD. The presence of artefacts away from the source would also play a role in the observed difference in the flux density.
In this work, we will discuss the spectro-temporal nature of the observed activity, discuss the physical mechanism and infer physical properties of the emission region in the corona.
3 Results
3.1 Results from VISAD
Figure 1a shows the Stokes V dynamic spectrum from VISAD made with 100 s and 10 MHz averaging in the top panel. The bottom panels shows the band averaged light curve obtained from VISAD in Stokes I and V, with 100 s averaging. Until 23:09:00 UT, we find multiple epochs of strong intermittent flaring activity, during which the stellar suddenly rises beyond 10 K with circular polarisation levels between 60 - 90%. Note that the values reported are averaged over the entire stellar disk. So the true will be higher depending on how small the area of the active region is compared to the size of the star. The flare rise phases are sudden and are followed by decay phases in both and polarisation fraction, as clearly seen in the band averaged light curves. We mark this initial 2 h period as a period of strong activity. Analysing the typical flux levels to which Stokes V emission falls back to after the sudden impulsive rise in the light curve, a basal Stokes V flux level is fixed at -2 mJy, corresponding to =-0.5 K. A similar exercise on the Stokes I light curve leads to a basal flux level of 4 mJy ( =1.0 K) which is about twice the typical quiescent flux seen during the quasi-steady period, as noted in Fig. 1. The basal flux levels are used to define flaring epochs during the strong activity period. A flaring epoch is said to begin when the stellar flux rises above the basal flux level and ends when the flux falls back below the basal level. Based on this definition, four flaring epochs (F1 - 4; see Fig. 1b-c) with flare decay timescales 5 - 30 min are identified in both of the Stokes light curves.
After the strong activity period, the star goes through 1.4 h period of weak activity between 2021-12-03 23:09:00 UT to 2021-12-04 00:24:00 UT (See, Fig. 1b-c). This period is characterised by high Stokes V 0.5 - 1.510, primarily the in higher frequency bands (700 MHz) as seen in VISAD spectrum (Fig. 1a). The observed Stokes V and Stokes I is much brighter than the expected coronal temperature of 3 MK based on X-ray studies (e.g., van den Besselaar et al., 2003; Namekata et al., 2020). Additionally, the band-averaged flux light curve in both Stokes I and V demonstrate high polarisation levels of 50 - 90% during the weak activity period, with a gradually varying profile resultant from the dynamic spectral structure of the emission. Such emission characteristics after flares have been observed in Sun as well and is interpreted as a signature of high-energy particles trapped in post-flare loops (Gergely, 1986; Gopalswamy, 2011; Hillaris et al., 2016; Morosan et al., 2019; Salas-Matamoros & Klein, 2020). We believe that the same phenomena is also behind the observed emission characteristics in this case as well.
After the weak activity period, we find that the Stokes V and I light curves tends to fluctuate around characteristic mean flux values of -0.5 and 1.5 mJy with relatively low circular polarisation levels. We define this phase as the quasi-steady phase.
3.2 Results from imaging analysis
To independently verify the flux density and polarisation trends presented in Section 3.1, we analysed the band-averaged light curve and image spectral cubes obtained from direct imaging (described in Section 2). Figure 2 shows the imaging-based light curve in Stokes I and V. Stokes V images have better SNR than those in Stokes I as evident from the lower error bars, due to relatively low levels of artefacts. Additionally, the number of galactic and extra-galactic sources with circular polarisation fraction exceeding are significantly low (e.g., Saikia & Salter, 1988; Macquart & Melrose, 2000; Enßlin et al., 2017; Callingham et al., 2023). Apart from the few time steps where only 5 upper limits could be determined either in V or I maps, we have good polarisation estimates across the light curve. As mentioned in Sect. 2, overall the fluxes from VISAD and imaging procedures agree within 20%. This flux uncertainty does not affect our analysis or physical modelling in Sect. 4. This is because, the emission being highly polarised in the flare and weak activity phase, is likely via some coherent emission processes. These coherent emission processes are often subject to a host of intrinsic physical processes, which are hard to nail down with just the available radio data (see, Bastian, 1990; Melrose, 2017b; Vedantham, 2021, for an overview). So our analysis in Sect. 4 requires absolute flux estimation accuracy only within an order of magnitude in (Bastian et al., 1990). The strong activity light curve matches well with the VISAD results and the significantly polarised weak activity period that followed is also well captured in the image-based light curve. During the weak activity period, Stokes I flux peaks at 2.5 mJy with a polarisation 60% similar to that seen in the VISAD light curve (Fig. 1(c)). A gradual decline phase after 00:40:00 UT is also seen in the Stokes I and V light curves (Fig. 2a-b), akin to the VISAD light curves (Fig. 1(b-c)).
In order to understand the spectral evolution of stellar emission, the stellar spectra derived from image spectral cubes described in Sect. 2 were analysed. Figure 3 shows example Stokes V spectra from specific time steps during the strong and weak activity period. Polarisation fraction is colour-coded. The spectrum shown in Fig. 3 for the strong activity period is chosen to coincide the peak epoch of the strongest flare (F3). The Stokes V flux of the star rises to around 16 mJy ( K) at high frequencies during this timestep (2021-02-03 22:12:13 UT) as seen in VISAD dynamic spectrum. Circular polarisation reaches to 100% during this time. The spectrum during the weak activity period shows an enhanced brightening with a significantly enhanced circular polarisation in the high frequency bands (700 MHz). This agrees with the observed enhanced confined to high frequency ranges in the VISAD dynamic spectrum during the weak activity phase (see, Fig. 1).
4 Discussion
The high and polarisation fraction seen in the strong activity period suggest a coherent emission mechanism at play, either via an ECME mechanism or a plasma emission process (see, Melrose, 1993, for an overview). The following subsections will summarise our current understanding of the two mechanisms.
4.1 Plasma emission
Plasma emission mechanism is the other coherent emission process which can also generate highly polarised emission (Tsytovich & Kaplan, 1969; Melrose & Sy, 1972). However, this mechanism is often subject to near-source scattering (e.g., Steinberg et al., 1971; Robinson, 1983; Arzner & Magun, 1999; Mohan et al., 2019; Kontar et al., 2019; Chen et al., 2020), which reduces the observed polarisation levels. In the case of solar type-III bursts, which are well known plasma emission sources, the maximum Stokes V/I detected is 60% (Saint-Hilaire et al., 2013; Rahman et al., 2020). However, these results are at frequencies below 100 MHz. Mohan (2021) showed that the effects of scattering could be less significant at higher frequencies since the effective scattering screen widths decreases towards higher frequency or equivalently at lower heights above corona. This means that the scattering induced depolarisation at our observing frequencies could be lower than that observed close to 100 MHz, but its effects may not be negligible. It is therefore unlikely that we observe radio bursts with polarisation levels above 75% generated purely via a coherent plasma emission mechanism.
4.2 Electron Cyclotron Maser Emission (ECME)
An ECME is driven by a loss-cone distribution of electrons. Such an unstable electron pitch-angle distribution can form in post-flare coronal loops (e.g. Wu & Lee, 1979; Winglee & Dulk, 1986; Ning et al., 2021). The ECME emission mechanism predicts that in /¡0.3 X-mode (polarised in the sense of electron gyration) emission is dominant at fundamental . Recent studies using fully kinetic particle-in-cell simulations suggest that a harmonic X-mode emission via Z-mode coalescence can also form at /¡0.3 and be observable (Ning et al., 2021). Meanwhile, in the 0.3¡/¡1.3 range, Z mode emission at dominates (Melrose et al., 1984b; Winglee, 1985) with polarisation in the opposite sense of electron gyration (O-mode). But, X-mode emission at higher harmonics of can also form by the coalescence of Z-mode waves when 0.3¡/¡1.3 regime (Melrose, 1991). In a narrow window of 1.3¡/¡1.5 X-mode emission emitted directly at the harmonics of can be observed (Melrose et al., 1984a). X-mode emission at harmonics of are expected via either a direct harmonic emission or various coalescence interactions up to /¡3 (Bastian et al., 1990). Harmonics higher than three are hardly produced via any of these process due to the requirement of very high energy density in the unstable wave modes (Melrose, 1991). So 3 is the highest observable harmonic expected. Another crucial factor is also that the pure X-mode emission at , or an X-mode emission at harmonics of , formed by Z-mode coalescence can be nearly 100% polarised (Melrose, 2009, 2017b).
When /¿ 3, the emissivity of the various ECME processes drops dramatically and along with the attenuation due to gyromagnetic absorption, ECME is expected to be insignificant (Melrose & Dulk, 1982; Bastian, 1990; Melrose, 1991; Ning et al., 2021).
To summarise, the possible X-mode polarised ECME pathways driven by loss-cone electron distribution are:
1.
/¡0.3: Fundamental emission at . Z-wave coalescence could produce emission observable at s (s = 2,3).
2.
0.3¡/¡1.3: Z/W wave coalescence can produce X-mode emission at s.
3.
1.3¡/¡3: Harmonic X-mode emission at s via either a direct emission or wave coalescence interactions.
Beyond /=3, plasma emission mechanism dominates. Ni et al. (2020) showed that when /, the Z, W and UH modes triggered by an electron cyclotron maser instability would not escape the plasma. Instead these waves can interact via a plasma emission mechanism to generate coherent emission at the fundamental and harmonic of the upper hybrid frequency, .
Besides, the existing theory suggests that O-mode emission can be produced via the following pathway:
1.
0.3¡/¡1.3: Z-mode pathway generates emission at .
4.3 The emission mechanism of AD Leo flares
In this subsection, we will discuss the various possible mechanisms for the observed emission given the possible local physical scenarios, during the various activity periods. First, we build a 1D density and magnetic field model for the star, using which and will be estimated across coronal heights. Zeeman Doppler Imaging (ZDI) based studies (Morin et al., 2008; Bellotti et al., 2023) infer that the magnetic field of AD Leo is predominantly dipolar () with surface field strengths within 0.1 - 0.5 kG. However, at strongly flaring active regions magnetic field strengths on AD Leo can be as high as 3kG (e.g., Cranmer & Saar, 2011). Based on this, we choose a typical range for field strength at the base of the active region, , between 0.3 and 1 kG. The chosen range considers a mean ZDI field strength at the lower end and the order of magnitude of the strongest reported active region —B— at the higher end. For the 1D analytical model of the radial magnetic field, two models are considered. First one is a dipole-dominated active region magnetic field model suggested for solar active regions by Aschwanden et al. (1999),
with the dipole scale height, assumed to be 0.1R (R is the stellar radius), based on solar studies (Aschwanden et al., 1999). Note that r is measured radially from stellar centre in units of R. In the second model, we assume a simple multipole expansion model of magnetic field with the net magnetic flux assumed to be distributed as 72% dipole, 21% quadrupole and 7% octapole fraction, based on the ZDI results for observations done in 2020 (Bellotti et al., 2023).
The fractional composition of the magnetic moments employed here, are based on large scale ZDI results that may not be representative of flaring active regions. Since there exist no estimation of multipole moments at the stellar active region, we assume the large-scale fractional distribution. However, the high relative dipolar concentration in Eqn. 2 assures that this is probably not an uncanny assumption for an active region magnetic structure if the AD Leo active regions are also similar to those on the sun with strong bipolar magnetic field concentrations and arcade-like structures. The two models will be referred to as Model 1 (Eqn. 1) and Model 2 (Eqn. 2) hereafter.
We derive assuming the density model () proposed by Villadsen & Hallinan (2019),
The basal coronal density in the model is derived from OVII line ratios by Ness et al. (2004). X-ray observations of pre-flare loops also suggest a basal coronal density within 10 - 10 (Namekata et al., 2020). Also, the assumed radial density profile is empirical, but consistent with the best-fit hydrostatic equilibrium profiles derived for solar active regions, based on direct measurements (e.g. Mann et al., 1999; Aschwanden et al., 2001). Zucca et al. (2014) showed that the data from a CME driving active solar region is well modelled by an exponential hydrostatic equilibrium density profile for r, which is the height range of interest in our study.
Figure 4(a) shows the derived and . The plotted range of correspond to 0.31 kG. The observing band (550 - 850 MHz) is marked in red. The green band shows the frequencies corresponding to one third of the observation band. First, consider Model 1 for the observing band. For an ECME emission process, the observing frequency, = s where s =1,2,3. The region where the red band overlaps the blue corresponds to the r range where =. Similarly, the r range where the green band overlaps the blue identifies the r range where =3. Beyond this r range, the possibility of ECME to be observed in 500 - 850 MHz is very low since inherent emissivity drops significantly (Melrose et al., 1984b). Vertical solid red lines mark the extent of the overlapping r range for =, and the solid green lines mark the r range for =3. Same exercise is performed for Model 2. Instead of solid lines, we use dotted lines of the same colour to mark the r ranges.
Figure 4(b) shows the ratio / for = , and /3. Horizontal lines mark /=0.3 ,1.3 and 3. We will first consider Model 1. The overlap between the red and blue bands give us the / values for the r range from where fundamental ECME mode emission can originate in the observing band. Clearly this ratio is greater than 1.3 but less than 3, implying that the fundamental X-mode emission is impossible but a Z-mode ECME is possible. However, the polarisation will be in the opposite sense of electron gyration (O-mode). However, an X-mode polarised emission is possible via the coalescence of Z or W-waves, as per the standard theory (Melrose, 1991). But, the resultant harmonic X-mode emission would not be observable in our band, but in 2 band. Meanwhile, the overlap between green and blue bands give us the / values in the r range from where if third harmonic emission arises, it can be observed in our . The range of / is much greater than 3, making ECME emission mechanism unfavourable. Similar to the above discussion, for Model 2, the intersection between gray and red band gives the range of / for fundamental ECME emission and the overlap between gray and green band gives the range of the ratio for harmonic emission to be observable in band. Consider the gray and red band overlap. We have cases for 0.3¡/¡3. Either a O or Z mode emission (0.3¡/¡1.3) or a harmonic X-mode (1.3¡/¡3) can be triggered here. But, akin to the case of Model 1, the harmonic X-mode emission born out of Z/W-mode coalescence or a direct emission cannot be observed in our observing band. However, the overlap between green (= /3) and gray band falls within 1.3¡/¡3 favouring a harmonic X-mode ECME from its r range via Z or W-mode coalescence or a direct harmonic emission, observable in band.
To summarise the discussion:
1.
Model 1 only favours an O-mode polarised (opposite sense of electron gyration) emission at via Z-mode pathway, observable in the band
2.
Model 2 favours an X-mode (polarised in the sense of electron gyration) emission at harmonics of local formed via either a direct emission or Z/W-mode coalescence pathways, and observable in band.
We now consider the polarisation in the bursts to further constrain the mechanism. The polarisation has consistently been left circular with Stokes V/I ¿ 60% during bursts. At higher frequencies (700 MHz), we find higher polarisation degrees attained during strong flares (e.g., Fig. 3). Morin et al. (2008) had shown that the south pole of AD Leo is inclined well along the line of sight (LOS) at an angle of 20. So there is a high chance that in the LOS magnetic field structures could be oriented southwards. This means the observed emission is in the sense of electron gyration,which means the emission is in X-mode. Also, the degree of polarisation is during flares, often reaching 100% at higher frequencies (see Fig. 3). If the emission is via an ECME process, the possible pathways are either a fundamental X-mode ECME mechanism from regions where = or a harmonic X-mode emission from regions where =/s and 0.3¡/¡3 (s=2,3). Based on the inferences from Fig. 4, Model 1 does not favour both fundamental or harmonic X-mode ECME, but Model 2 favours a harmonic X-mode emission by all possible pathways. Meanwhile, Zhang et al. (2023) observed the star in the 1 - 1.5 GHz band during the strong activity period, between 21:13 – 22:48 UT, and reported emission features that are better explained by an ECME mechanism. Unfortunately their observations did not extend further into the weak and quiescent activity phases.
However, this analysis does not explicitly consider the propagation effects that would affect the observed ECME. Without multi-waveband spectroscopic data, the existing stellar atmospheric models (e.g. Fontenla et al., 2016; Wedemeyer et al., 2013) cannot be well constrained to derive a reliable ambient/quiescent physical structure (especially, B(r) and ), crucial to assess the atmospheric opacity (see White & Kundu, 1997; Treumann, 2006, for an overview). Consequently, based on current knowledge on the ECME mechanism as derived fromsolar observations, the emissivity should fall drastically above (see, Sect. 4.2). Moreover, the emission for higher harmonics (s=) is less affected by gyromagnetic absorption from upper coronal layers (Zheleznyakov, 1970; Melrose & Dulk, 1982). Both these strongly favour the detection of ECME at s=2,3. Our inferences on ECME generation based on 1D stellar models also favour harmonic emission.
Plasma emission, which is the competing mechanism to ECME produces coherent emission observable at (F-branch) and 2 (H-branch) (Ginzburg & Zhelezniakov, 1958; Tsytovich & Kaplan, 1969; Melrose & Sy, 1972; Melrose, 2009). This mechanism generally favours polarisation in the O-mode (Melrose, 2017b) especially for the F branch. For the H branch, this is true if , but if /1.5, highly polarised X-mode emission at =2 is possible (Willes & Melrose, 1997). Given that we observe a high X-mode polarised flares, it is likely a H-branch emission from regions with /1.5. From Fig. 5b, we infer that the H-band emission observable in should arise from r1.9R, where ¡/2. In this r range, ¿¿ for Model 1. But for some B values, /¡1.5 is satisfied for Model 2. So only Model 2 can support a H-band plasma emission in X-mode.
Based on the discussions so far, Model 2 favours X-mode polarised emission via both H-branch plasma emission mechanism and ECME emission processes at harmonics of . Note that Model 1 does not support any of the likely emission processes favouring a strong X-mode ECME or H-band plasma emission. This possibly means that a solar like active region model is not good for M dwarfs, but a ZDI based multipole expansion model works better. For the rest of the discussion, we will hence consider only Model 2. Meanwhile, the H-branch plasma emission observed in the sun has always shown polarisation levels below 30% (Willes & Melrose, 1997; Saint-Hilaire et al., 2013; Rahman et al., 2020). So, during strong activity period, given the high polarisation of 60 - 100% (Figs. 1-3) the emission mechanism is likely via an ECME mechanism.
In the analysis so far, we had kept the basal coronal density constant in the model. Let us consider a variation of this value within an order of magnitude and assess the possible ECME pathways supported by the two B(r) models. For this analysis, we vary n within 10 - 10 cm, consistent with the values reported by previous spectroscopic and X-ray continuum observations. This corresponds to a variation in the assumed basal coronal density in Eq. (3) by 0.4 to 4 times. If the basal density goes up 4 times, since we would see a an increase in by a factor of 2, which in turn raises all the / bands in Fig. 4 by the same factor. Note that in Fig. 4 the green band overlaps with gray (Model 2) within 1¡/¡3. So the harmonic X-mode emission pathway, previously supported by Model 2 will remain feasible but now in the 2¡/¡3. The other ECME pathways that were already not supported by the previously assumed lower density plasma will anyway be unsupported by a denser plasma due to enhanced damping of the ECME driving instability. However, since /3 overlaps with Model 2 in /¡10, an ECME driven coherent plasma emission process suggested by Ni et al. (2020) can produce harmonic X-mode emission observable in . Meanwhile, if we assume a 0.4 times lower density in the active region, all the / bands in Fig. 4 will be reduced by a factor of 1.6. In this case for Model 2, harmonic X-mode emission will be observable as earlier. Interestingly, a portion of the overlapping region between the Model 1 (blue band) the green band will fall within 2.5¡/¡3, suggesting a possibility of observing a harmonic X-mode emission in . This may make Model 1 viable for a small range of B and basal density values. However, Model 2 consistently supports an ECME mechanism for an order of magnitude wide range of B and n values.
4.4 Weak activity period - type-IV burst
The post-flare weak activity phase is found to last for 1.4 h. The Stokes V dynamic spectrum (Fig. 1a) shows that the post-flare emission shows a bright patch confined to frequencies 700 MHz, between 2021-12-03 23:27 and 2021-12-04 02:25 UT. To highlight this better a 200 s averaged VISAD spectrum was made (see, Fig. 6a). The imaging-based spectrum from the weak activity period (Fig. 3b) also shows a characteristic high polarisation bright emission feature above 700 - 730 MHz, where the emission polarisation jumps above 55% and rises to 80% as frequency increases. As per the solar radio burst classification scheme, the aforementioned features suggest that the weak activity phase corresponds to a type-IV radio burst confined to frequencies above 700 MHz (Gopalswamy, 2011; Hillaris et al., 2016; Salas-Matamoros & Klein, 2020). For comparison, Fig. 6b shows a solar dynamic spectrum during a strong CME taken by Wind/WAVES instrument. The solar data is taken from the publicly available type-IV catalogue compiled by Coordinated Data Analysis Workshops (CDAW) data center 333https://cdaw.gsfc.nasa.gov/CME_list/radio/type4. The solar event was so chosen to match the dynamic spectral emission profile of the AD Leo. The morphological analogues of the type-III and type-IV bursts in the solar template spectrum can be identified in the AD Leo DS.
Type-IV radio bursts are associated with strong flares and CMEs in the Sun, like the much sought-after type-II bursts (Robinson, 1986; Gopalswamy, 2011; Hillaris et al., 2016; Kumari et al., 2021). While a type-II burst is associated with the coronal shock generated by a fast CME, a type-IV burst is generated by flare-accelerated electrons trapped either in the post-flare loops or in the moving plasmoids associated with the CME (e.g., Wild, 1970; Smerd, 1970; Gopalswamy et al., 2005; White, 2007; Gopalswamy, 2011). Type-IVs are generally associated with relatively longer and stronger solar X-ray flares than type-IIs (Cane & Reames, 1988). Also, type-IV bursts are often linked to strong CMEs which expand out to heliocentric angular widths 60, with mean speeds 900 km s (Robinson, 1986; Gopalswamy, 2011; Hillaris et al., 2016). Recently, Zic et al. (2020) reported the first type-IV burst detection in a star other than the Sun, namely Proxima Centauri which is an old (5 Gyr) M dwarf. Given its age ( 250 Myr) and rotation period (P = 2.2 days), AD Leo belongs to a fast rotating (period 5 days), young (age 1 Gyr) and more active stellar population (‘C’ branch) compared to the the Sun (age = 4.6 Gyr; P = 25 days), which belongs to the less active ‘I’ branch stars (Barnes, 2003; Brown, 2014). Barnes (2003) demonstrated that the stars undergo an evolutionary migration from ‘C’ to ‘I’ branch as they spin down with age due to mass and angular momentum loss. This migration involves a fundamental change in the way the internal dynamo couples with the external active atmospheric layers and the stellar wind (Brown, 2014; Garraffo et al., 2018). So, though Proxima Centauri is an M dwarf, it is an ‘I’ branch star, but AD Leo is not. Here, we report the first detection of a type-IV burst on a young (250 Myr), active ‘C’ branch M dwarf.
We do not find evidence for a type-II burst, making it still elusive in an M dwarf despite several day-long monitoring campaigns on various stars, including AD Leo (e.g., Osten et al., 2008; Villadsen & Hallinan, 2019). Though rare, solar type-II bursts have also been found associated with coronal shocks produced by failed eruptions, unrelated to CMEs (e.g., Magdalenić et al., 2012; Morosan et al., 2023). AD Leo is an active star with an order of magnitude stronger and predominantly axi-symmetric magnetic field (Moschou et al., 2019). Theoretical models have argued that the strong magnetic field structures and the associated high Alfvén speed barriers in the stellar corona could either block even the energetic flares from causing an eruption/CME (failed eruption scenario), or resist the formation of a strong shock required to initiate a type-II emission during an eruption/CME (Odert et al., 2020; Alvarado-Gómez et al., 2022). This means that, unlike the case of the sun, AD Leo could host more failed eruptions during strong flares, and even if a CME/eruption occurs, type-II bursts need not be triggered for most events. These scenarios can explain why have there been no report of a type-II burst on AD Leo despite several monitoring campaigns (e.g., Osten & Bastian, 2006; Osten et al., 2008; Villadsen & Hallinan, 2019). In any case, since post-flare loop structures will trap flare-accelerated electrons, strong type-IV bursts could trace CMEs better than type-IIs, in M dwarfs.
Given the high X-mode polarisation level and high at 700 MHz, and based on our inferences for the active region in the previous section, the type-IV burst emission likely originates from an ECME mechanism in the Z/W mode coalescence or a harmonic emission pathways. An ECME mechanism is usually triggered by an unstable loss-cone distribution of energetic electrons formed within a magnetic bottle like configuration, readily provided by the post-flare loops (e.g., Carley et al., 2019). In the case of the sun, type-IV bursts are generally associated with plasma emission mechanism in the metric waveband (Weiss, 1963; Salas-Matamoros & Klein, 2020). However, during some strong flares, associated type-IV bursts have demonstrated ms to s scale spiky emission features in their dynamic spectra with high degree of circular polarisation and , often modelled as generated by an ECME mechanism (e.g., Carley et al., 2019; Liu et al., 2018; Morosan et al., 2019; Cliver et al., 2022). Certain high time resolution studies of AD Leo in 1-3 GHz observations have reported ms scale spiky emission features (e.g., Osten et al., 2008; Zhang et al., 2023). The time resolution of uGMRT data restricts us from exploring sub-second scale fine structures in the dynamic spectrum. The ECME signatures we detect in the type-IV emission above 700 MHz could probably be formed by numerous sub-second scale fine structures. Meanwhile, at frequencies below 700 MHz the emission polarisation is well below 30%. The emission mechanism here is likely a H-branch plasma emission which is known to produce high emission at low polarisation levels (e.g., Willes & Melrose, 1997; Saint-Hilaire et al., 2013; Rahman et al., 2019).
Since the type-IV emission is seen to be mostly confined within 700 MHz (Fig. 6), using Model 2 and assuming that the ECME is formed at either the second or the third harmonic of , the height of the post flare loop can be inferred. Consider Fig. 5. It can be seen from Fig. 5b that within 1.8 - 2.25 R, the overlap region of cyan and gray bands satisfy the condition of 0.3¡/¡1.5 required for Z-mode/harmonic X-mode mechanism to operate efficiently and produce a second harmonic ECME observable in . Since 700 MHz is roughly the centre of cyan band, the post-flare loop height could be 1.8 - 2.0 R. Similarly, if the emission is at the third harmonic (= 3), we should use Fig. 4(a-b) to obtain emission height. Here instead of the cyan band, we consider the overlap between the green and the gray bands. The emission height range turns out to be 2.0 -2.5 R, implying a post-flare loop height of 2 - 2.25 R. This implies that the star can have loops rising to about 1R above the photosphere.
4.5 Quasi-steady period
The quiescent Stokes I flux level varies within 1 - 1.75 mJy (see Figs. 1c & 2) implying values of the order of 10 K. However, the polarisation levels of 20 - 30% inferred from the light curve suggests a H-band coherent plasma emission mechanism. The observed emission is likely a superposition of various spatially separated coherent plasma emission regions or a mixture of gyrosynchroton and plasma emission regions (Bastian et al., 1990; Bastian, 1994).
5 Conclusions
We present a two hour-long metric band flaring activity in a young active M dwarf, AD Leo. A hydrostatic equilibrium electron density profile was assumed with a basal density ranging within 10 - 10 cm. Based on the existing knowledge on the stellar magnetic field structure, different radial field profiles (B(r)) were considered: a solar-like model and a simple multipole expansion model. Contrasting the radio data with these models we conclude the following:
Active region B(r) is more consistent with a multipole expansion model than a solar-like model, given the possible range of B(r) and basal coronal density.
The observed radio bursts are powered by an ECME mechanism.
The period of strong flaring was succeeded by a post-flare activity period. We report a solar like type-IV radio burst in AD Leo, in the post-flare phase characterised by a distinctive dynamic spectral morphology and circular polarisation levels ranging within 60 - 100%. Previously, a type-IV emission was reported in Proxima Centauri by Zic et al. (2020). This is the first solar-like type-IV detection on a young active M dwarf belonging to a different age-activity population (‘C’ branch) relative to the Sun (belonging to ‘I’ branch) (Barnes, 2003). The emission characteristics favour an ECME mechanism at the second or third harmonic of the local gyrofrequency. The detection of the type-IV burst could signify a strong CME, since solar type-IVs are known to be associated with fast and wide CMEs. Nevertheless, an hour-long type-IV burst signifies powerful accelerated particle activity in the post-flare loops. Based on the frequency extent of the type-IV burst in the VISAD dynamic spectrum we infer that the post-flare loops could rise up to 1 R scales above photosphere into the inter-planetary space. The quasi-steady emission that followed the type-IV activity is likely powered by a mixture of several coherent plasma emission regions.
Acknowledgements
This work is supported by the Research Council of Norway through the EMISSA project (project number 286853) and the Centres of Excellence scheme, project number 262622 (“Rosseland Centre for Solar Physics”). We thank the staff of the GMRT that made these observations possible. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. This research made use of NASA’s Astrophysics Data System (ADS). AM and NG are supported in part by NASA’s STEREO project and LWS program. AM acknowledges the developers of the various Python modules namely Numpy (Harris et al., 2020), Astropy (Astropy Collaboration et al., 2013), Matplotlib (Hunter, 2007) and multiprocessing. AM also thanks the developers of CASA (McMullin et al., 2007). AM acknowledges the discussions with Barnali Das which greatly helped speed up the data analysis. AM acknowledges the discussions with Tim Bastian.